please help me understand what to do with the following question.
4x-12x+27=0

Answer by ewatrrr(24785)

 
Hi,
Solving by doing allowable operations
I. Combine Like terms
II. Isolate 'unknown x' on one side of the equation
III. read the value of x
4x-12x+27=0
-8x = -27 |subtracting 27 from both sides of EQ
x = 27/8 or 3 3/8